OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -6 and a vertex at (-8, -8)?

OpenStudy (anonymous):

x = -10 x +10 = 0 x=-6 x+6 = 0 (x+10)(x+6)=0

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you are welcome remember to multiply that out